User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



VSN: HAY-0050 (+1025 GB)

  • Exp started OK (JS).
  • Perpetual alarm from error m5 -104 mk5cn: time-out, connection closed. The field system and markV still responds to commands thouygh (JS)
  • weather monitor not working, going to get Brett to reboot it tomorrow (Imogen)
  • 11:20 UT - 70K stage is at 100K with 20K stage at 65K (Ross).


VSN: OSOD-038 (+0 GB)

  • Exp started OK (JS).
  • 06:16 UT - Large clock drift detected during check, watching for stability (Ross).
  • 06:26 UT - Clock clearly unstable, performed fmset in between scans for now (Ross).
  • 06:32 UT - dbbc reset performed in between scans. First safe scan is 300-0634a (Ross).
  • 10:06 UT - Clock starting to drift, few dbbc restarts and fmsets to get stable, missing scans from 300-1007b onwards. (Ross).
  • 10:53 UT - Field system has become unresponsive and recording cannot be resumed. Field system restarted and recording resumed with scan 300-1116 (Ross). – Missing 60 GB


VSN: NRCAN-09 (+0 GB)

  • Exp started OK (JS).
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1711.txt · Last modified: 2015/10/27 11:23 by Ross Turner